# references.bib

### v1

@@ 2.1.4 Necessary Conditions for E2EE Confidentiality
End-to-end encryption provides its promised confidentiality guarantee only under certain conditions: namely, when (1) the underlying encryption method is secure and (2) that encryption method is securely implemented in the application being used.
@@
Even if the math underlying an encryption scheme is perfect in theory, the process of deploying the encryption scheme in software or hardware introduces avenues for security failures, and the use of the encryption scheme by fallible humans in practice introduces further potential for security failures.
In order for E2EE confidentiality to hold: (1) the code implementing the E2EE application must be correct, (2) the hardware device that the application runs on must be secure, and (3) the users using the application must use it correctly and take adequate security precautions. [^11] In fact, these are essential considerations for the proper functioning of any software product, not limited to encryption. Undermining any one of these three could undermine E2EE confidentiality: for example, (1) the software might fail to perform the encryption correctly; (2) the hardware device might fail to generate a secret key correctly, or it might be infected by malware; and (3) a user might fail to keep their secret key secret. Errors can be introduced in myriad ways: through simple mischance, human mistake, or intentional sabotage.
@@ 2.2 Background on AI Assistants
AI assistants are programs designed to interpret everyday language and perform computational tasks. Today’s AI assistants, such as ChatGPT, Gemini, and Llama, are able to handle a wide range of tasks, including text analysis, content creation, code generation, language translation, and more. At the core of these technologies are statistical models—programs trained on data to identify patterns. Specifically, AI assistants typically rely on large language models (LLMs), which are trained on vast datasets of text to generate outputs based on predictions of the next word. Due to the size of their training datasets and sophisticated architectures, LLMs can handle complex and previously unseen user inputs (queries) and provide contextually relevant responses.
@@ 2.2.1 Training and Inference
AI assistants undergo two key phases: training, when the model is built to identify patterns, relationships, and rules; and inference, when real-time responses are generated based on user inputs. Training itself often involves multiple stages, typically referred to as pretraining and fine-tuning. In pretraining, the model takes in vast, general datasets to learn structures and patterns in the data. Fine-tuning then refines this pretrained model by training it on additional, more specialized datasets, improving its accuracy and relevance in particular contexts. We describe these processes in further detail below.
@@
Models have configurable parameters which are numerical values that determine how outputs are generated. These parameters are used in calculations to predict outputs when the model is queried with an input. During training, the model processes large datasets—such as books, websites, code repositories, and conversational datasets in the case of LLMs—to adjust (or “tune”) its parameters. This tuning happens over many iterations of training, during which the model’s parameters capture more and more refined patterns in the data. The finalized set of parameters constitutes the trained model, which can then be prompted with inputs (e.g., questions) to generate outputs (e.g., answers). Models at the scale of commercial AI assistants contain billions or even trillions of parameters trained over the course of months on high-performance computers.
@@
Once trained, models can perform inference, during which they generate real-time responses to user inputs. When a user inputs a query to the model, it is processed through many steps using the model parameters to calculate an output. While inference is less resource-intensive than training, it still requires significant computational power and is typically handled by centralized servers.
Deployed models undergo continuous improvement through updates that may involve additional training on more diverse or recent datasets, reinforcement learning with human feedback, where collected user feedback is used to improve response quality; patches to address issues such as specific problematic outputs, and model refinement with new techniques and architectures. As such, inference and training often overlap in a feedback loop, where data collected during inference can be fed back into the training pipeline to fine-tune the model further. This process is shown in Figure 2.
Figure 2: Inference and training continuously happen with an AI assistant as a feedback loop. As users continue to query the AI assistant, this generates data that is stored and can be used to then train and improve the model.
@@ 2.2.2 Data Collection
Once the AI assistant is deployed, various types of data may be collected. This can include chat logs, which record user inputs (text, voice, or image data) and the system’s responses. These historical records, even if kept by the platform, may or may not be visible to users. [^12] Search/query history can be surprisingly revealing when aggregated over time, involving family concerns, health issues, religious beliefs, and more. Research has shown users can be reluctant to disclose all their search history even for better personalized search results [PanjwaniSSJ13]. Additionally, metadata, such as timestamps, device type, language, and geographic region, as well as user preferences and settings, may also be collected and stored by the platform. The collection of this data serves multiple purposes, which can be broadly categorized as follows:
@@
    Ease of use: The data is used to enable certain features, such as maintaining conversation histories, or maintaining continuity within a session or across sessions, for an individual user.
@@
These categories are not mutually exclusive, as data may be collected to serve any or all of these purposes. The collection of user data is commonly enabled by default with collected user data transmitted and stored on centralized cloud servers. In some cases, some or all user data may remain local on the user’s device, eliminating external transmission.
@@ 2.2.3 Architecture of AI Assistants in Messaging Platforms
Next, we overview the architecture of AI assistants integrated in messaging platforms. The architecture is how the tool is structured and has implications on how and where data is processed. First, it is important to distinguish between the messaging application, the AI assistant, and the model. The messaging application is the application installed on the user’s device which facilitates communication between users and provides the interface for interacting with the AI assistant, e.g., WhatsApp, Messenger, and Signal. The AI assistant is the system that performs tasks such as answering queries, generating responses, or providing recommendations. At its core is the model—in this case, an LLM—the underlying computational engine that processes inputs and generates outputs based on patterns learned during training.
@@
Here we differentiate between two types of architectures: cloud-based and local, illustrated in Figure 3.
@@
2.2.4 Availability in E2EE Applications
AI assistants are already available in major E2EE applications as a standalone tools outside of conversations between users. For example, WhatsApp features a search bar that can be used both to search through conversations and to “Ask Meta AI”. When the user inputs a search query, the results show matches from their conversation history (via a local search) as well as an option to invoke Meta AI with the search query. In this case, we do not describe Meta AI as being integrated with E2EE content, as only explicit Meta AI queries initiated by the user are processed by Meta AI, without any E2EE content. The availability of such non-integrated features in E2EE applications is outside the scope of our concern in this paper, which is focused on AI models processing E2EE content.
In contrast, integrating AI assistants as a feature within E2EE conversations would typically require the AI assistant to process E2EE content. For example, this could be in the form of allowing users to invoke the AI assistant within an ongoing conversation to provide contextually relevant support or responses.
@@ 2.3 How Is It Technically Possible for E2EE Data to Be Fed Into AI Assistants?
The current discourse around using E2EE data in AI assistants, therefore, is focused primarily on the idea of feeding readable plaintext versions of E2EE messaging data into AI assistants. Since only the sender and recipient devices in an E2EE system can read the plaintext messages, the sender or recipient devices would be the only ones who could provide the plaintext data to the AI assistants. This possibility appears to be primary focus of the ongoing discourse: that users’ devices would be set up to send E2EE content to cloud-based AI assistants for processing.
Nothing about E2EE technically prevents this from happening—indeed, no technical measure would both be able to allow senders and recipients to read their own messages, and prevent them from showing them to anyone else. This is much like no matter how secretively you encode and deliver your hand-written letter personally to your friend, nothing you can do can stop your friend from showing the decoded letter to other people once they have it.
@@ 2.4 Background on Trusted Execution Environments (TEEs)
Trusted hardware refers to physical devices that are designed to ensure the confidentiality and integrity of computations, even if the owner of the hardware may be compromised. Trusted hardware is designed to create a secure area of computation, in which no information about the computation is released outside of the trusted hardware during execution of code, and that code will run exactly as specified. There are many types of technologies that fall under the broad umbrella of trusted hardware—such as hardware security modules (HSMs), trusted platform modules (TPMs), and trusted execution environments (TEEs)—which offer different capabilities. Of these technologies, TEEs are the most relevant for the implementation of AI functionalities, as they are designed to run arbitrary programs; other technologies currently have more limited, bespoke use cases, such as key generation and storage. We thus focus on TEEs in this subsection, and in the rest of the paper.
We can conceputalize a TEE as a physical “lockbox”, which is designed to execute arbitrary code such that no outside entities (including, for example, the operating system, other device processes, and third parties) can observe nor manipulate the state of the code as it is running. During initial setup of the TEE, a computer program—call it $\Pi$—is loaded inside the TEE. The TEE then has a single input/output interface, through it which receives inputs $x$ from outside the TEE, computes the program $\Pi$ on the inputs, and outputs the result $\Pi(x)$ from the TEE. During execution of $\Pi$, the TEE maintains some internal ephemeral state in memory, which stores temporary values during computation. TEEs are also often equipped with persistent storage, on which they can store data longer-term. See  [pass2017formal] for a formal treatment of TEE abstractions.
@@
The goal of TEE design is to enable outsourcing of computations over user data to potentially untrusted (but more computationally powerful) environments, simply by loading the program inside a server-side TEE, and processing user data inside of it. As described so far, the inputs to and outputs from the TEE can be seen by the owner of the TEE. However, inputs and outputs can be hidden from the TEE owner with an additional step: encrypt inputs and outputs so that they can only be decrypted using secret key that is stored inside the TEE. Then, computations inside the TEE can still happen in plaintext, within the “shield” of the hardware protections provided by the TEE.
@@ 2.4.1 Necessary conditions for TEE security
The security guarantees of TEEs naturally rely on the correct implementation of the TEE, and that the hardware has not been tampered with at any point in the TEE’s life cycle (e.g., tamper-proof seals have not been removed, backdoors have not been inserted, etc). This requires trust in various actors across the hardware supply chain, from manufacturing, to assembly, to deployment. Further, users need assurance that the application provider is actually running a legitimate TEE on their servers. While this can be implemented via the TEE’s attestation capabilities, the attestation service needs to leverage some trusted authority, such as the hardware provider.
Even if the hardware is correctly implemented according to specification, the security of TEEs additionally relies on the specification itself being free of vulnerabilities. Unfortunately, there is a history of practical attacks on various TEE implementations [van2024sok, fei2021security]. These attacks range across a variety of threat models—including adversaries that have physical access to the TEE, can launch co-located TEEs with arbitrary programs, control the system’s boot process, etc—and the adversary’s goal is to violate the confidentiality and integrity of the TEE, i.e., recover sensitive runtime state or tamper with the data in the TEE. Notable examples of attack vectors include physical and micro-architectural side-channels, which exploit shared caches [brasser2017software], branch predictors [lee2017inferring], page faults [xu2015controlled], transient execution [van2020lvi], and more. This history of attacks has prevented TEEs from reaching mainstream adoption for security applications.
@@ 2.4.2 TEEs vs E2EE
The security of E2EE and TEEs rely on two broad assumption types: (1) specification correctness, and (2) implementation correctness. The former states that the design of the technology/protocol is secure, while the latter states that a particular instance of the technology/protocol is correctly implemented according to the specification. TEEs and E2EE have important differences for both assumption types. Assumption type (1) generally hinges on the hardness of a mathematical problem for E2EE; and on a robust hardware design for TEEs. On the other hand, assumption type (2) hinges on the security of software and hardware implementations for E2EE and TEEs, respectively. These entail a completely different set of considerations between E2EE and TEEs, since the software and hardware supply chains involve different entities, vendors, and processes.
@@
The difference in the type of security offered by TEEs and by encryption mean that depending on the context, one may be more suitable for use than the other. That said, when considering whether to use TEEs or encryption for a given application that might admit either, the former raise certain kinds of security concerns not present for the latter. Widely used encryption schemes today have largely withstood decades of scrutiny without critical confidentiality breaks; indeed, new encryption schemes that have not yet been subject to such scrutiny are widely considered unsuitable for use in security applications. Our understanding and scrutiny of TEE security is comparatively young, and the field has seen and continues to see a stream of discoveries of security flaws with impacts on deployed systems (e.g., [van2024sok, fei2021security]); at the same time, TEEs enable secure computation in promising new applications, which would not be practical based on cryptography alone.
2.5 A Practical Example: Apple Intelligence

### v2

@@ 2.1.4 Necessary Conditions for E2EE Confidentiality
End-to-end encryption provides its promised confidentiality guarantee only under certain conditions: namely, when (A) the underlying encryption method is secure and (B) that encryption method is securely implemented in the application being used.
@@
Even if the math underlying an encryption scheme is perfect in theory, the process of deploying the encryption scheme in software or hardware introduces avenues for security failures, and the use of the encryption scheme by fallible humans in practice introduces further potential for security failures. In order for E2EE confidentiality to hold: (1) the code implementing the E2EE application must be correct, (2) the hardware device that the application runs on must be secure, and (3) the users using the application must use it correctly and take adequate security precautions. [^9] In fact, these are essential considerations for the proper functioning of any software product, not limited to encryption. Undermining any one of these three could undermine E2EE confidentiality: for example, (1) the software fails to perform the encryption correctly; (2) the hardware device is infected with malware; or (3) a user fails to keep their secret key secret. Errors can be introduced in myriad ways: through simple mischance, human mistake, or intentional sabotage.
@@
2.1.5 E2EE Messaging Features and Usability
Many academic definitions of E2EE messaging consider a simple model where the only activity that occurs is literal messaging—sending messages between senders and receivers and within group chats. However, modern messaging apps are more feature-rich, and the ways in which such additional features process message contents—sometimes in tension with E2EE confidentiality—has long been a subject of debate. Some common features are, strictly speaking, in violation of E2EE confidentiality: for example, link previewing, where the app displays a “preview” of a link contained in a message, technically violates E2EE confidentiality because a third party processes the link in order to fetch the preview. Furthermore, some common features of E2EE applications enhance confidentiality despite not being part of the basic definition of E2EE: for example, user blocking helps protect users from harassment and spam, but has no direct relationship to E2EE confidentiality (or any of the other cryptographic security properties mentioned in Section 2.1.2). [^10] Other common features with potential consequences for confidentiality and E2EE confidentiality under some implementations include: importing contacts, autocorrect, and accessibility features such as text-to-voice.
In short, any comprehensive treatment of E2EE messaging applications must consider the interaction of commonplace messaging features with the system. As discussed in Section 2.1.3, an E2EE application requires a sufficiently large and diverse user base to promise privacy in practice, and it is therefore pragmatically important for E2EE messaging apps to incorporate common usability features of messaging apps if doing so could make or break users’ willingness to choose the app. Of course, this is not an argument for a kitchen-sink approach to messaging app development: some features may be too incompatible with E2EE, or can at least be made more palatable by choosing certain implementations over others. For example, Signal has link previewing, but also offers users the ability to disable link previewing in the settings.
This complicated relationship between messaging features, usability, and privacy in E2EE apps has led to practical controversy as well as academic debate. For example, Apple offers a feature where users can store encrypted backups of their iPhone data on Apple’s cloud storage—a standard convenience. In February 2025, reports emerged that the UK government has ordered Apple to grant it backdoor access to these encrypted backups [apple-uk-backups]. In this case, Apple providing (encrypted) backups—providing utility to the end user—also created a new resource for third parties (here, law enforcement) to target to gain access to end-to-end encrypted user communications, potentially without users’ knowledge and without breaking encryption.
While it is out of our scope to fully characterize the wide spectrum of messaging app features and their privacy implications, it is useful to view AI assistants in the context of a broader constellation of other privacy-impacting messaging app features which have been met with controversy and debate.
@@ 2.2 Background on AI Assistants
AI assistants are programs designed to interpret everyday language and perform computational tasks. Today’s AI assistants, such as ChatGPT, Gemini, and Llama, are able to handle a wide range of tasks, including text analysis, content creation, code generation, language translation, and more. At the core, these technologies are statistical models—programs trained on data to identify patterns. Specifically, AI assistants are typically built using large language models (LLMs), which are trained on vast datasets of text to generate outputs based on predictions of the next word. Due to the size of their training datasets and sophisticated architectures, LLMs can handle complex, previously unseen user inputs (queries) and provide contextually relevant responses.
@@ 2.2.1 Training and Inference
AI assistants undergo two key phases: training, when the model is built to identify patterns, relationships, and rules; and inference, when real-time responses are generated based on user inputs. Training itself often involves multiple stages, typically referred to as pretraining and fine-tuning. In pretraining, the model takes in large-scale general datasets to learn structures and patterns in the data. Fine-tuning then refines this pretrained model by training it on additional, more specialized datasets, improving its accuracy and relevance in particular contexts. We describe these processes in further detail below.
@@
Models have configurable parameters which are numerical values that determine how outputs are generated. These parameters are used in calculations to predict outputs when the model is queried with an input. During training, the model processes large datasets—such as books, websites, code repositories, and conversational datasets in the case of LLMs—which iteratively adjusts (or “tune”) the parameters, capturing more and more refined patterns in the data with repeated passes. The finalized set of parameters constitutes the trained model, which can then be prompted with inputs (e.g., questions) to generate outputs (e.g., answers). Models at the scale of commercial AI assistants contain billions or even trillions of parameters, trained over the course of months on high-performance computers.
@@
Once trained, the model can be used for inference, where it outputs real-time responses to user inputs. When a query is inputted to a model, it is processed with the model parameters to calculate an output. While inference is less resource-intensive than training, it still requires significant computational power and is typically handled by centralized servers.
Deployed models undergo continuous improvement through updates that often involve additional training on more diverse or recent datasets. Some approaches for this include reinforcement learning with human feedback, where collected user feedback is used to improve response quality [ouyang2022training]; patches to address issues such as specific problematic outputs, and model refinement with new techniques and architectures. As such, inference and training often overlap in a feedback loop, where data collected during inference is fed back into the training pipeline to fine-tune the model further. This process is shown in Figure 2.
Figure 2: Inference and training of AI assistants happen continuously as a feedback loop. As users continue to query the AI assistant, this generates data that is stored and can be used to further train and improve the model.
@@ 2.2.2 Data Collection
Once the AI assistant is deployed, various types of data from the user interactions with the assistant may be collected. This can include chat logs, which record user inputs (text, voice, or image data) and the system’s responses. These historical records, even if kept by the platform, may or may not be visible to users. [^11] Search/query history can be surprisingly revealing when aggregated over time, involving family concerns, health issues, religious beliefs, and more. Research has shown that users can be reluctant to disclose all their search history even for better personalized search results [PanjwaniSSJ13]. Additionally, metadata, such as timestamps, device type, language, and geographic region, as well as user preferences and settings, may also be collected and stored by the platform. This data serves multiple purposes, which can be broadly categorized as follows.
@@
    Ease of use: The data is used to enable certain features, such as maintaining conversation histories or continuity within a session/across sessions, for an individual user.
@@
Collecting user data is commonly enabled by default with collected user data transmitted and stored on centralized cloud servers. In some cases, some or all user data may remain local on the user’s device, eliminating external transmission.
@@ 2.2.3 Architecture of AI Assistants in Messaging Platforms
Next, we overview the architecture of AI assistants integrated in messaging platforms. The architecture is how the application is structured technically, which has implications on how and where data is processed. First, it is important to distinguish between the messaging application, the AI assistant, and the model. The messaging application is the application installed on the user’s device which facilitates communication between users and provides the interface for interacting with the AI assistant, e.g., WhatsApp, Messenger, and Signal. The AI assistant is the system that performs tasks such as answering queries, generating responses, or providing recommendations, e.g., Meta AI, which is built into WhatsApp and Messenger. At its core is the model—in this case, an LLM—the underlying computational engine that processes inputs and generates outputs based on patterns learned during training.
@@
We differentiate between two types of architectures: cloud-based and local, illustrated in Figure 3.
@@
2.2.4 Standalone vs. Integrated AI Features in E2EE Applications
AI assistants are commonly available in major E2EE applications, as a standalone tool separate from user conversations (and any other E2EE content). For example, an E2EE messaging application might also include a search bar where users can type in queries to search the web and/or send to AI models.
In general, we do not consider such standalone tools to be integrated with E2EE applications in the way we are concerned about in this paper. Standalone AI features separated from E2EE content generally do not raise the security concerns that we argue do arise with integrated AI features, unless providers create the impression that these AI features are themselves E2EE. We consider standalone features out of our scope. (Section 4.3 provides further discussion with practical examples based on Meta AI features in WhatsApp.)
Our focus is, rather, on integrated AI models and their capacity to process E2EE content. Integrating AI assistants as a feature within E2EE conversations would often require the AI assistant to process E2EE content: for example, to allow users to invoke the AI assistant within an E2EE conversation or to provide contextually relevant support, responses, or summaries within E2EE conversations.
@@ 2.3 How Is It Technically Possible for E2EE Data to Be Fed Into AI Assistants?
The current discourse around using E2EE data in AI assistants, therefore, is focused primarily on the idea of feeding readable plaintext versions of E2EE messaging data into AI assistants. Since only the sender and recipient devices in an E2EE system can read the plaintext messages, the sender or recipient devices would be the only ones who could provide the plaintext data to the AI assistants. This possibility appears to be primary focus of the ongoing discourse: that users’ devices would be set up to send plaintext copies of E2EE content to cloud-based AI assistants for processing.
Nothing about E2EE technically prevents this from happening—indeed, no technical measure would both be able to allow senders and recipients to read their own messages, and prevent them from showing them to anyone else. This is a bit like no matter how secretively you encode and deliver your hand-written letter personally to your friend, nothing you can do can stop your friend from showing the decoded letter to other people once they have it. (This analogy falls short after a point. In the letter-writing example, your friend would likely have to take an affirmative, intentional action to show the letter to others; but in the encrypted messaging case, users’ devices and applications can be set up by service providers to automatically send plaintext copies of messages to third parties without any action being taken by users.)
@@ 2.4 Background on Trusted Execution Environments (TEEs)
Trusted hardware refers to physical devices that are designed to ensure the confidentiality and integrity of computations, even if the owner of the hardware may be compromised. Trusted hardware is designed to create a secure area of computation, in which no information about the computation is released outside of the trusted hardware during execution of code, and that code will run exactly as specified. There are many types of technologies that fall under the broad umbrella of trusted hardware—such as hardware security modules (HSMs), trusted platform modules (TPMs), and trusted execution environments (TEEs)—which offer different capabilities. Of these technologies, TEEs are the most relevant for the implementation of AI functionalities, as they are designed to run arbitrary programs; other technologies currently have more limited, bespoke use cases, such as key generation and storage. We thus focus on TEEs from here on.
We can conceptualize a TEE as a physical “lockbox”, which is designed to execute arbitrary code such that no outside entities (including, for example, the operating system, other device processes, and third parties) can observe nor manipulate the state of the code as it is running. During initial setup of the TEE, a computer program—call it $\Pi$—is loaded inside the TEE. The TEE then has a single input/output interface, through it which receives inputs $x$ from outside the TEE, computes the program $\Pi$ on the inputs, and outputs the result $\Pi(x)$ from the TEE. During execution of $\Pi$, the TEE maintains some internal ephemeral state in memory, which stores temporary values during computation. TEEs are also often equipped with persistent storage, on which they can store data longer-term. See [pass2017formal] for a formal treatment of TEE abstractions.
@@
The goal of TEE design is to enable outsourcing of computations over user data to potentially untrusted (but more computationally powerful) environments, by loading the program inside a server-side TEE, and processing user data inside of it. As described so far, the inputs to and outputs from the TEE can be seen by the owner of the TEE. However, inputs and outputs can be hidden from the TEE owner with an additional step: encrypt inputs and outputs so that they can only be decrypted using secret key that is stored inside the TEE. Then, computations inside the TEE can still happen in plaintext, within the “shield” of the hardware protections provided by the TEE.
@@ 2.4.1 Necessary conditions for TEE security
The security guarantees of TEEs naturally rely on the correct implementation of the TEE, and that the hardware has not been tampered with at any point in the TEE’s life cycle (e.g., tamper-proof seals have not been removed, backdoors have not been inserted, etc). This requires trust in various actors across the hardware supply chain, from manufacturing, to assembly, to deployment. Further, users typically do not handle TEEs directly; rather, they are embedded in devices or handled server-side. As such, users rely on the application or device provider that a TEE is in fact being used for any given application.. While some TEEs have additional mechanisms to attest to their usage for certain computations (e.g., via their attestation service), this generally needs to leverage some trusted authority, such as the hardware provider.
Even if the hardware is correctly implemented according to specification, the security of TEEs additionally relies on the specification itself being free of vulnerabilities. Unfortunately, there is a history of practical attacks on various TEE implementations [van2024sok, fei2021security]. These attacks range across a variety of threat models—including adversaries that have physical access to the TEE, can launch co-located TEEs with arbitrary programs, control the system’s boot process, etc.—and violate the confidentiality and integrity of the TEE, i.e., recover sensitive runtime state or tamper with the data in the TEE. Some of the attacks have been mitigated since their discovery; others have not, and cannot be fully mitigated with known techniques [van2024sok]. Notable examples of attack vectors include physical and micro-architectural side-channels, which exploit shared caches [brasser2017software], branch predictors [lee2017inferring], page faults [xu2015controlled], and transient execution [van2020lvi].
@@ 2.4.2 TEEs vs E2EE
The security of E2EE and TEEs rely on two broad assumption types: (1) specification correctness, and (2) implementation correctness. The former states that the design of the technology/protocol is secure, while the latter states that a particular instance of the technology/protocol is correctly implemented according to the specification. TEEs and E2EE have important differences for both assumption types. Assumption type (1) generally hinges on a computational hardness assumption for E2EE; and on a robust hardware design for TEEs. On the other hand, assumption type (2) hinges on the security of software and hardware implementations for E2EE and TEEs, respectively. These entail a completely different set of considerations between E2EE and TEEs, since the software and hardware supply chains involve different entities, vendors, and processes.
@@
The difference in the type of security offered by TEEs and by encryption mean that depending on the context, one may be more suitable for use than the other. That said, when considering whether to use TEEs or encryption for a given application that might admit either, the former raise certain kinds of security concerns not present for the latter. Widely used encryption schemes today have largely withstood decades of scrutiny without critical confidentiality breaks; indeed, new encryption schemes that have not yet been subject to such scrutiny are widely considered unsuitable for use in security applications. Our understanding and scrutiny of TEE security is comparatively young, and the field has seen and continues to see a stream of discoveries of security flaws with impacts on deployed systems (e.g. [van2024sok, fei2021security]); at the same time, TEEs enable secure computation in promising new applications, which would not be practical based on cryptography alone.
2.5 Deployments in Practice: Apple, Samsung, and Meta AI
In this section, we overview three deployed examples of AI assistants that may process E2EE data: Apple Intelligence, Samsung’s Galaxy AI, and Meta AI in WhatsApp. We apply our recommendations to the case of Apple Intelligence in Section 7.3, as we found that the documentation currently available on Apple Intelligence supports a more detailed analysis than for the other two products. We leave the other two examples—as well as future releases of these and other products—as an exercise for the reader.
The information in this section is based on publicly available documentation from the relevant vendors, not direct testing of the products. We list key points that currently available documentation leaves unclear, under the heading “What We Still Don’t Know” at the end of each subsection. These lists are not comprehensive, but highlight notable points of ambiguity that hinder a complete evaluation of the E2EE compatibility of these AI integrations.
The latest versions of two of the products that we cover here were launched since the last version of this paper was published, less than three months ago. So, while our case-specific description and analysis will remain illustrative, some details may soon be out of date. Readers seeking to understand the implications of our work for the latest technologies should check recent product releases and may conduct their own analysis based on our generalizable analytical framework and recommendations.
2.5.1 Apple Intelligence